OpenStudy (anonymous):

Please help! :) 49^ 5x+2 = (1/7)^ 11-x

umerlodhi (umerlodhi):

1 sec

OpenStudy (anonymous):

\[49=7^2\\ \frac{1}{7}=7^{-1}\] is a start

OpenStudy (anonymous):

Yea i thought of 49= 7^2 but idk where to put the 2 in 5x+2

OpenStudy (triciaal):